OpenStudy (anonymous):

may you help me with these questions please.answers only :) Question 34 The graph of which quadratic equation has a vertex of (3, 5)? A) y = (x – 3)2 + 5 B) y = (x + 5)2 + 3 C) y = (x + 3)2 + 5 y = (x – 5)2 + 3 Question 36 Solve. 2x2 + x = 15. A) x = 2 and x = –15 B) x = 17 and x = 13 C) x = and x = –3 D) x = –2 and x = Question 38 When solving x2 + 6x = 10 by the completing the square method, what value is added to both sides? A) 4 B) 9 C) 3 D) 5 Question 39 Solve. x2 + 8x + 25 = 0 A) –4 ± 3i B) –4 ± i C) –3 ± 4i D)

OpenStudy (anonymous):

Question 39 Solve. x2 + 8x + 25 = 0 A) –4 ± 3i

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

Question 38 When solving x2 + 6x = 10 by the completing the square method, what value is added to both sides? 6x/2=3x 3x^2=9x B) 9

OpenStudy (anonymous):

ohh okay i see how you got it.thank you :)

OpenStudy (anonymous):

By y = a(x - h)² + k form where (h,k) is the vertex, we can show that.. "y = (x – 3)^2 + 5" is the equation with that vertex

OpenStudy (anonymous):

A) y = (x – 3)2 + 5

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

but whats #36?